OpenStudy (anonymous):

In a small bag of candy, there are 9 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(You select first an orange candy and then a blue candy).

OpenStudy (anonymous):

probability first is orange is?

OpenStudy (lexipoo1998):

What is P suppose to represent?

OpenStudy (anonymous):

In a small bag of candy, there are 9 orange candies and 4 blue candies. You select 1 candy, eat it, and then select another. Find P(You select first an orange candy and then a blue candy). A. 9/26 B. 36/169 C. 1/36 D. 3/13

OpenStudy (anonymous):

the probability the first is orange is \(\frac{9}{13}\)

OpenStudy (anonymous):

then the probability that the second is blue, if the first was orange, is \(\frac{4}{12}=\frac{1}{3}\)

OpenStudy (anonymous):

multiply these two numbers together for your answer

OpenStudy (anonymous):

is it 3/13